OpenStudy (randomuser):

Show your work in solving the equation. Include the work to check your solution and show that your solution is extraneous

OpenStudy (randomuser):

I need help on how to do an extraneous equation on radicals

OpenStudy (isaiahhlindsayy):

what is the question you need help on\

OpenStudy (randomuser):

I need help on this one I already got my numbers . Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a√x+b+c=d

OpenStudy (randomuser):

I do not know how to make an extraneous problem

OpenStudy (isaiahhlindsayy):

An extraneous solution is a root of a transformed equation that is not a root of the original equation because it was excluded from the domain of the original equation. Example 1: Solve for xx , 1x − 2+1x + 2=4(x − 2)(x + 2)1x − 2+1x + 2=4(x − 2)(x + 2) . 1x − 2+1x + 2=4(x − 2)(x + 2)1x − 2+1x + 2=4(x − 2)(x + 2) (x − 2)(x + 2)(x − 2)+(x − 2)(x + 2)(x + 2)=4(x − 2)(x + 2)(x − 2)(x + 2)(x − 2)(x + 2)(x − 2)+(x − 2)(x + 2)(x + 2)=4(x − 2)(x + 2)(x − 2)(x + 2) (x−2)+(x+2)=4(x−2)+(x+2)=4 2x=42x=4 x=2x=2 But 22 is excluded from the domain of the original equation because it would make the denominator of one of the fractions zero--and division by zero is not allowed! . Therefore, it cannot be a root of the original equation. So, 22 is an extraneous solution. So, the equation has no solutions.

OpenStudy (randomuser):

Ok but how would I do it with these numbers I do not want you to tell me just guide me they are - 2 6 4 8 What order should I put them in I already tried this once and got this part wrong.

OpenStudy (randomuser):

It is in task 2

OpenStudy (isaiahhlindsayy):

Step 1 : Pulling out like terms : 1.1 Pull out like factors : 2x + 2 = 2 • (x + 1) Equation at the end of step 1 : Step 2 : Equations which are never true : 2.1 Solve : 2 = 0 This equation has no solution. A a non-zero constant never equals zero. Solving a Single Variable Equation : 2.2 Solve : x+1 = 0 Subtract 1 from both sides of the equation : x = -1 x = -1

OpenStudy (isaiahhlindsayy):

kind of like this?

OpenStudy (randomuser):

My teacher wants with radicals and I am not good with radicals

OpenStudy (isaiahhlindsayy):

@skullpatrol

OpenStudy (randomuser):

@3mar

OpenStudy (3mar):

Well, I am here. I was praying!

OpenStudy (randomuser):

ok cool could you help me on task 2 on the extraneous problem

OpenStudy (3mar):

Ok. But if you are not in hurry, it would take some time!

OpenStudy (randomuser):

im not

OpenStudy (3mar):

That is good!

OpenStudy (randomuser):

doyou know my numbers?

OpenStudy (3mar):

" my numbers?"???

OpenStudy (3mar):

Ok. What you got for part 1?

OpenStudy (randomuser):

That part was just to show the numbers I picked

OpenStudy (randomuser):

they are 2, 4, 6, 8 how do I make an extraneous problem out of those numbers?

OpenStudy (randomuser):

@3mar

OpenStudy (3mar):

Time for pray - 15 min I will be back In Sha' Allah after 12-20 min

OpenStudy (randomuser):

are you there?

OpenStudy (3mar):

I am here now

OpenStudy (randomuser):

Ok could you help me?

OpenStudy (3mar):

After you did  what is the question?